# A Classification Theorem for Steady Euler Flows

### v1

@@
Abstract
Fix an analytic simply connected domain $\Omega\subset\mathbb{R}^{2}.$ A smooth function $\psi:\Omega\rightarrow\mathbb{R}$ that is constant on $\partial\Omega$ is said to define a steady Euler solution on $\Omega$ if the Poisson bracket between $\psi$ and its Laplacian vanishes. In other words, $\psi$ is a steady Euler solution if
$\{\psi,\Delta\psi\}=0.$ ($\star$)
[…]
1 Introduction
The study of steady Euler flows is old and vast, the first such solutions appearing in Euler’s original paper [19]. The interest in steady flows is natural, both from a mathematical and practical point of view. From the mathematical side, steady flows play a distinguished role in the long-time dynamics of all solutions as they do in any dynamical system. Indeed, the first step to understand the long-time dynamics of any system is to have a deep and complete understanding of the steady solutions. While it is well known that the Euler equation admits an infinite-dimensional family of steady states on any two-dimensional domain, little is known about the global properties of the set of solutions. The purpose of this work is to take a step in that direction.
1.1 The Euler equation
We begin by recalling the 2d Euler equation for incompressible flow on a simply connected domain $\Omega\subset\mathbb{R}^{2}$:
$\partial_{t}\omega+u\cdot\nabla\omega=0,$ (1.1)
$u=\nabla^{\perp}\Delta^{-1}\omega.$ (1.2)
@@ 1.1 The Euler equation
One of the most captivating aspects of the 2d Euler equation is in the long-time behavior of solutions. An amazing fact, borne out of numerical and physical experiments, is that 2d Euler solutions starting from “any old data” tend to relax to much simpler states in the long-time limit. This is conjectured to occur, despite the conservative nature of the equation and the absence of a clear relaxation mechanism. Recent breakthroughs have been achieved in our understanding of this in certain perturbative regimes starting with the work [4] and then the later works [27, 28, 35]. See also [15] for a recent work leveraging the classical stability theory of steady states to make some long-time statements. Understanding the mechanism for this type of behavior outside of the perturbative regime appears well outside of the reach of any of the available techniques in the field. In order to be able to study such large data phenomena, perhaps in the long-time limit, it is necessary to build a flexible theoretical framework that will allow us to study questions of a more global nature. For us, this begins with having a robust understanding of Euler steady states.
1.2 Steady States
@@
in which case the steady Euler equation can be written as:
$\nabla^{\perp}\psi\cdot\nabla\Delta\psi:=\{\psi,\Delta\psi\}=0,$ (1.3)
where we note that $\psi$ is constant on $\partial\Omega.$ Classically, there are three classes of solutions that have been considered:
@@
The presence of the first two is clearly due to the symmetries of the Laplacian and the Poisson bracket. The third type is based on the simple observation that
@@
These three classes of steady solutions have a very nice structure. A quite robust linear stability theory has been developed for the first two, through the study of the Rayleigh equation [13, 31, 41]. The stability theory for the third type can be investigated using properties of the linearized operator associated to (1.4):
@@
see [1, 2, 3, 32, 33]. Given the tools the community has developed over the years to study these three types of steady states, it would be quite fortunate if we were able to show that these are the only steady states out there. This, and more, has been established locally in some non-degenerate settings [9, 10, 11, 17] (here, non-degenerate means that the Schrödinger operator $\mathcal{L}$ above is invertible).
In recent years, there has appeared a quite large literature on rigidity results for steady states. These results can be classified into two types: those making structural assumptions and those making smallness assumptions. In a series of papers, Hamel and Nadirashvili [23, 24, 25] established a number of rigidity results on symmetric domains assuming, for example, that the velocity field does not vanish on the domain. For example, they proved in [25] that sufficiently regular steady states on a regular annulus, say $B_{2}(0)\setminus B_{1}(0)$, with non-vanishing velocity must be radial. Of course, as is exemplified by the Dirichlet eigenfunctions, the non-vanishing of the velocity field is essential to such a result. See also the work of Li, Lv, Shahgholian, and Xie [30] on rigidity of Poiseuille flows in a strip and the work of Gui, Xie, and Xu[22] that studies the rigidity of steady Euler flows in terms of the flow angles. Another interesting recent result, among others, given by Gómez-Serrano, Yao, Park, and Shi [21] establishes the radial symmetry of solutions with compactly supported and non-negative vorticity. On the side of rigidity results under smallness conditions, there are now a few results showing that any solution sufficiently close to some special shear flow must be a shear flow [11, 34]. As an example, Lin and Zeng [34] proved that traveling-wave Euler solutions that are sufficiently close to the Couette flow on $\mathbb{T}\times[0,1]$ must be steady shear flows. Finally, we mention the work of Constantin, Drivas, and Ginsberg [10] that establishes results on both flexibility and rigidity in general domains, the main question being whether stable solutions inherit the symmetries of the domain. See also the works [7, 8, 16, 37] for further results on flexibility and rigidity in various contexts. We now state our main result.
1.3 Main Result
@@
Let $\Omega$ be an analytic, simply connected, and bounded subset of $\mathbb{R}^{2}$. Assume that $\psi$ is analytic on $\bar{\Omega}$, is constant on $\partial\Omega$, and satisfies the steady Euler equation. Let $[a,b]$ be the range of $\psi$ on $\bar{\Omega}.$ Then, one of the following must hold.
@@
We show moreover that $F$ is actually Hölder continuous globally, admitting a convergent Puiseux series at the endpoints. In fact, the precise form of the Puiseux expansion of $F$ at $a$ and $b$, see Lemma 4.5, allows us to define $\mathcal{L}=\Delta-F^{\prime}(\psi)$ as a self-adjoint operator on suitable (weighted) spaces. In particular, it can be decomposed into an invertible Schrödinger operator plus a compact perturbation and the Fredholm alternative can be applied to study its invertibility properties. This is key to understanding the local structure of the set of steady states near arbitrary ones. We summarise this as an interesting consequence of the proof in the following corollary.
@@
Assume $\psi\in C^{\omega}(\bar{\Omega})$ for an analytic, simply connected, and bounded domain $\Omega.$ If
@@
and if $[a,b]$ is the range of $\psi$ on $\bar{\Omega},$ then $F$ is analytic on $(a,b)$ and admits a (locally) convergent Puiseux series at the endpoints.
@@
The main novelty of Theorem 1.1 compared to other rigidity results on steady Euler flows mentioned above is that we make no structural or smallness assumptions on $\Omega$ or on $\psi$. The results are also, in some sense, optimal since [18] shows that we cannot relax the analyticity condition to finite smoothness and Theorem 4.12 shows the result cannot hold in general on multiply connected domains. The analyticity assumption is one drawback of Theorem 1.1; it would be nice if it could be replaced by a less stringent structural assumption.
1.4 Main ideas
@@
We now move to discuss some of the main ideas that go into the proof of Theorem 1.1. As we have mentioned above, the proof of Theorem 1.1 is based on a number of observations about the Poisson bracket, the zero sets of analytic functions, and about over-determined elliptic problems. To clarify the main points of the argument, we will briefly sketch some of the main ideas here.
1.4.1 Some observations about $\{f,g\}=0$
It is well known, using a straightening lemma, that if $f,g\in C^{1}$ and $\{f,g\}=0$ in a neighborhood of some $x_{*}\in\mathbb{R}^{2}$, while $\nabla f(x_{*})\not=0,$ then we can write
$g=F(f),$
@@ 1.4.1 Some observations about $\{f,g\}=0$
for some $F\in C^{1}$ in a sufficiently small neighborhood of $x_{*}.$ This is a simple application of the inverse function theorem. It is also clear that this need not be the case globally. There are two important examples one should keep in mind in this regard, both of which we give on $\mathbb{T}^{2}$ (though the examples can be easily modified to work on the unit disk, for example).
@@
Example 1: Non-Analytic $g$. Take $f(x,y)=\sin(x)\sin(y).$ In this case, each regular level set of $f$ has two connected components consisting of simple closed curves. We may take $g$ to simply be a bump function that is constant along the level sets of $f$ restricted to the first quadrant and also compactly supported inside the first quadrant. In this case, $g$ cannot be expressed as a function of $f$ globally on $\mathbb{T}^{2}$ (since $f$ takes the same values in the first and third quadrants, but $g$ takes different values there). We give two illustrative figures as follows.
Figure 1: 2D Contour plot of $g$
Figure 2: 2D Contour plot of $f$
@@
Example 2: A wall of critical points. Take $f=\sin^{2}(x)\sin^{2}(y)$ and $g=\sin(x)\sin(y).$ Then, $g$ cannot be written as $F(f).$ In this case, $f$ and $g$ are both analytic; however, the set of critical points of $f$ partitions the domain into multiple cells surrounded by walls of critical points. We give two illustrative figures below. The red lines in the figure below are the set of critical points of $f$.
Figure 3: 2D contour plot of $g$
Figure 4: 2D contour plot of $f$
@@
Assume that $f,g$ are analytic functions on some open, connected, domain $\Omega.$ Assume that $\{f,g\}=0$ and that the critical points of $f$ of even order are isolated. Then, there exists a continuous function $F:\mathbb{R}\rightarrow\mathbb{R}$ so that
@@
1.4.2 Properties of the set $\{\nabla\psi=0\}$ when $\{\psi,\Delta\psi\}=0.$
Given the preceding discussion, the obstruction to having a global relation between $\Delta\psi$ and $\psi$ in a 2d Euler steady state is the existence of a closed curve of critical points of $\psi$ of even order. Since $\psi$ is analytic, we can consider the inner-most such curve, $\Gamma$. On the region enclosed by $\Gamma$, called $\Omega_{\Gamma}$, $\psi$ will satisfy a semi-linear elliptic equation with non-linearity $F$ and the full gradient of $\psi$ vanishes on $\Gamma$. Note that the curve $\Gamma$ may have singularities. Now there appear three cases:
@@
    $\psi$ vanishes higher than quadratically on $\Gamma$,
@@
    $\psi$ vanishes only quadratically (somewhere) on $\Gamma$ and the associated non-linearity $F$ is not analytic at $c=\psi(\Gamma)$,
@@
In case (iii), it turns out that $F$ may be extended to a larger region containing $\Gamma$ and eventually we return to cases (i) and (ii). In both cases (i) and (ii), because $F$ is necessarily non-analytic at $c=\psi(\Gamma)$, it turns out that $\psi-c$ is necessarily single signed in $\Omega_{\Gamma}$–this is a key aspect of the argument. In case (i), then we can apply a known result of Brock [5] to deduce that $\psi$ must be radial; a nice feature of Brock’s result [5] is that it does not require the boundary of the domain to be smooth. The analysis in case (ii) is more difficult and will be discussed in more detail below. We give an illustration depicting the potential complexity of the level-set in the figure below [^2] The solid curves on the figure above are the set of critical points of $\psi$ in an elliptical domain. The boundary of region A demonstrates situation (i), where all the points are critical points of degree higher than 2. The boundary of region B demonstrates the second case (ii), in the case where all of the points are quadratic. Regions C and D could represent the third case, where $F$ could be extended from C to D, while D falls into the case (i) or (ii)..
Figure 5: Illustrations for a potential critical set of $\psi$
One snag in case (ii) is that we have to show that this curve $\Gamma$ is actually smooth and that $\psi$ is single signed in its interior. For general analytic functions, it is well known that the zero set may be quite complicated and may contain multiple singular points of different types [29, 6], even more so in high dimension. Note, for example, that the set of critical points in Example 2 above has a singularity at the origin. In the figure above, we have to show that case (ii) above actually corresponds to a region like B in the figure (i.e. a region with a smooth boundary). A key observation here is that, when we have that $\{\psi,\Delta\psi\}=0,$ the degree of vanishing of $\nabla\psi$ on each connected component of $\Gamma$ will necessarily be constant. This is established by using the constancy of $\Delta\psi$ on the level sets of $\psi$. Once the constancy of the degree is established, we use an elementary fact: singularities in the zero set of an analytic function occur only at points where the degree of vanishing of the function changes. The constancy of the degree of vanishing thus implies that $\Gamma$ is smooth.
1.4.3 Moving Plane Arguments
From the preceding two steps, we see that if a steady state $\psi$ does not satisfy the global relation $\Delta\psi=F(\psi)$, up to subtracting a constant, we end up with a simply connected domain $\Omega_{\Gamma}$ with analytic boundary where

### v2

@@
Abstract.
Fix an analytic simply connected domain $\Omega\subset\mathbb{R}^{2}.$ We show that all analytic steady states of the Euler equations with stream function $\psi$ are either radial or solve a semi-linear elliptic equation of the form $\Delta\psi=F(\psi)$ with Dirichlet boundary conditions. In particular, if $\Omega$ is not a ball, then there exists a one to one correspondence between analytic steady states of the Euler equations and analytic solutions of equations of the form $\Delta\psi=F(\psi)$ with Dirichlet boundary conditions.
Key words and phrases:
Steady Euler equations, stream function, analytic, critical point, moving plane method
1991 Mathematics Subject Classification:
35Q31, 35N25, 35J15, 35B53
1. Introduction and Main Results
The study of steady Euler flows is old and vast, the first such solutions appearing in Euler’s original paper [24]. The interest in steady flows is natural, both from a mathematical and a practical point of view. From the mathematical side, steady flows play a distinguished role in the long-time dynamics of all solutions, as they do in any dynamical system. Indeed, the first step to understanding the long-time dynamics of any system is to have a deep and complete understanding of the steady solutions. While it is well known that the Euler equations admit an infinite-dimensional family of steady states on any two-dimensional domain, little is known about the global properties of the set of solutions. The purpose of this work is to take a step in that direction.
1.1. The Euler Equations
We begin by recalling the 2D Euler equations for incompressible flow on a simply connected bounded domain $\Omega\subset\mathbb{R}^{2}$:
(1.1) $\partial_{t}\omega+u\cdot\nabla\omega=0,$
(1.2) $u=\nabla^{\perp}\Delta^{-1}\omega.$
@@ 1.1. The Euler Equations
One of the most captivating aspects of the 2D Euler equation is in the long-time behavior of solutions. An amazing fact, borne out of numerical and physical experiments, is that 2D Euler solutions starting from “any old data” tend to relax to much simpler states in the long-time limit. This is conjectured to occur, despite the conservative nature of the equation and the absence of a clear relaxation mechanism. Recent breakthroughs have been achieved in our understanding of this in certain perturbative regimes, starting with the work [4] and then the later works [32, 33, 40]. See also [19] for a recent work leveraging the classical stability theory of steady states to make some long-time statements. Understanding the mechanism for this type of behavior outside the perturbative regime appears to be well outside the reach of any of the available techniques in the field. In order to be able to study such large data phenomena, perhaps in the long-time limit, it is necessary to build a flexible theoretical framework that will allow us to study questions of a more global nature. For us, this begins with having a robust understanding of Euler steady states.
1.2. Steady States
@@
in which case the steady Euler equations can be written as:
(1.3) $\nabla^{\perp}\psi\cdot\nabla\Delta\psi:=\{\psi,\Delta\psi\}=0,$
where we note that $\psi$ is chosen to be zero on $\partial\Omega.$ Classically, there are three classes of solutions that have been considered:
@@
The presence of the first two is due to the symmetries of the Laplacian and the Poisson bracket. The third type is based on the simple observation that
@@
These three classes of steady solutions have a very nice structure. A quite robust linear stability theory has been developed for the first two, through the study of the Rayleigh equation [18, 37, 48]. The stability theory for the third type can be investigated using properties of the linearized operator associated with (1.4):
@@
see [3, 2, 1, 38, 39]. Given the tools the community has developed over the years to study these three types of steady states, it would be great if we were able to show that these are the only steady states out there. This, and more, has been established locally in some non-degenerate settings [12, 13, 15, 22] (here, non-degenerate means that the Schrödinger operator $\mathcal{L}$ above is invertible).
In recent years, there has appeared a quite large literature on rigidity results for steady states. These results can be classified into two types: those making structural assumptions and those making smallness assumptions. In a series of papers, Hamel and Nadirashvili [28, 29, 30] established a number of rigidity results on symmetric domains assuming, for example, that the velocity field does not vanish on the domain. For example, they proved in [30] that sufficiently regular steady states on a regular annulus, say $B_{2}(0)\setminus B_{1}(0)$, with non-vanishing velocity must be radial. Of course, as is exemplified by the Dirichlet eigenfunctions, the non-vanishing of the velocity field is essential to such a result. See also the work of Li, Lv, Shahgholian, and Xie [35] on the rigidity of Poiseuille flows in a strip and the work of Gui, Xie, and Xu[27] that studies the rigidity of steady Euler flows in terms of flow angles. Another interesting recent result, among others, given by Gómez-Serrano, Park, Shi, and Yao [25] establishes the radial symmetry of solutions with compactly supported and non-negative vorticity. On the side of the results of rigidity under smallness conditions, there are now some results showing that any solution sufficiently close to some special shear flow must be a shear flow [15, 36]. As an example, Lin and Zeng [36] proved that traveling-wave Euler solutions that are sufficiently close to the Couette flow on $\mathbb{T}\times[0,1]$ must be steady shear flows. Finally, we mention the work of Constantin, Drivas, and Ginsberg [13] that establishes results on both flexibility and rigidity in general domains, the main question being whether stable solutions inherit the symmetries of the domain. See also the works [8, 9, 21, 42] for further results on flexibility and rigidity in various contexts. We now state our main results.
1.3. Main Results
@@
Let $\Omega$ be a simply connected and bounded domain of $\mathbb{R}^{2}$. Assume that $\psi$ is analytic on $\bar{\Omega}$, is constant on $\partial\Omega$, and satisfies the steady Euler equation (1.3). Let $[a,b]$ be the range of $\psi$ on $\bar{\Omega}.$ Then, one of the following must hold.
@@
As $\psi$ is analytic in $\overline{\Omega}$ and $\partial\Omega$ is a level set of $\psi$, due to the structure theorem on the zero sets of real analytic functions (See Proposition 2.3), the boundary $\partial\Omega$ has to be a piecewise analytic curve.
Remark 1.2.
We show moreover that $F$ is actually Hölder continuous globally, admitting a convergent Puiseux series at the endpoints. In fact, the precise form of the Puiseux expansion of $F$ at $a$ and $b$, see Lemma 4.4, allows us to define $\mathcal{L}=\Delta-F^{\prime}(\psi)$ as a self-adjoint operator on suitable (weighted) spaces. In particular, it can be decomposed into an invertible Schrödinger operator plus a relatively compact perturbation. Hence the Fredholm theory can be applied to study its invertibility properties. This is key to understanding the local structure of the set of steady states near arbitrary ones. We summarize this as an interesting consequence of the proof in the following corollary.
@@
For a given analytic, simply connected and bounded domain $\Omega$, Suppose that $\psi\in C^{\omega}(\bar{\Omega})$ (the set of real-analytic functions in $\bar{\Omega}$) takes a constant on $\partial\Omega$. If
@@
and if $[a,b]$ is the range of $\psi$ in $\bar{\Omega},$ then $F$ is analytic on $(a,b)$ and admits a (locally) convergent Puiseux series at the endpoints.
@@
The main novelty of Theorem 1.1, compared with other results of rigidity in steady Euler flows mentioned above, is that we make no structural or smallness assumptions on $\Omega$ or on $\psi$. The results are also, in some sense, optimal, since [23] shows that we cannot relax the analyticity condition to finite smoothness, and Proposition 4.11 shows that the result cannot hold in general on multiply connected domains.
Remark 1.3.
The analyticity of the velocity field is the major assumption of Theorem 1.1. It would be nice if one could give a necessary and sufficient condition for the regularity of the velocity field so that Theorem 1.1 holds. On the other hand, for any $s>1$, there exists a single-variable compactly supported function in the Gevrey class $G^{s}$ (see [34]). Hence, we can use two compactly supported disjoint radial bumps to construct a steady flow whose stream function belongs to the Gevrey class $G^{s}$ with $s>1$, but is neither radial nor a solution to a global semilinear elliptic equation. Therefore, in terms of regularity, analyticity might be the optimal assumption for velocity fields.
Remark 1.4.
The simple connectedness of the domain is necessary. In Section 4.3, a counterexample is provided to illustrate that in a multiply connected domain, there exists a flow that can be neither described by a solution of global semilinear equation for the stream function nor a radial flow.
Remark 1.5.
As an application of Theorem 1.1, we can show that in a simply connected convex bounded domain, any analytic steady state whose vorticity is single-signed should inherit the mirror symmetry of the domain. In particular, any analytic steady flow with distinguished-signed vorticity in the disk is a radial flow. For more details on this property, one may refer to the Appendix B.
Remark 1.6.
To better understand the long-time behavior of unsteady flow, it would be desirable to establish a similar result in a broader class of steady states. An important fact that could help us to get closer is that every steady state with $\psi\in C^{2,\alpha}$ has analytic Lagrangian trajectories [10, 45, 14]. In fact, this can be extended to all steady states with merely bounded vorticity. The main idea is that steady Euler solutions with bounded vorticity actually have some stratified regularity away from the critical points, as in the case of regular vortex patches [44, 11, 5]. We will present this result in a forthcoming paper.
1.4. Main Ideas
@@
We now move on to discuss some of the main ideas that go into the proof of Theorem 1.1. As we have mentioned above, the proof of Theorem 1.1 is based on a number of observations about the Poisson bracket, the zero sets of analytic functions, and over-determined elliptic problems. To clarify the main points of the argument, we will briefly sketch some of the main ideas here.
1.4.1. Some observations about $\{f,g\}=0$
@@
It is well known, using a straightening lemma, that if $f,g\in C^{1}$ and $\{f,g\}=0$ in a neighborhood of some $x_{*}\in\mathbb{R}^{2}$, while $\nabla f(x_{*})\not=0,$ then we can write $g=F(f)$ for some $F\in C^{1}$ in a sufficiently small neighborhood of $x_{*}.$ This is a simple application of the inverse function theorem. It is also clear that this need not be the case globally. There are two important examples one should keep in mind in this regard, both of which we give in $\mathbb{T}^{2}$ (although the examples can be easily modified to work on the unit disk, for example).
@@
Example 1: Non-Analytic $g$. Take $f(x,y)=\sin(x)\sin(y).$ In this case, each regular level set of $f$ has two connected components consisting of simple closed curves. We may take $g$ to be a bump function that is constant along the level sets of $f$ restricted to the first quadrant and also compactly supported inside the first quadrant. In this case, $g$ cannot be expressed as a function of $f$ globally on $\mathbb{T}^{2}$ (since $f$ takes the same values in the first and third quadrants, but $g$ takes different values there). We give two illustrative figures as in Figure 1.
Figure 1. 2D Contour plot of $g$ and $f$ in Example 1
@@
Example 2: A wall of critical points. Take $f=\sin^{2}(x)\sin^{2}(y)$ and $g=\sin(x)\sin(y),$ then $g$ cannot be written as $F(f).$ In this case, $f$ and $g$ are both analytic; however, the set of critical points of $f$ partitions the domain into multiple cells surrounded by walls of critical points. We give two illustrative figures in Figure 2, where the red lines are the set of critical points of $f$.
Figure 2. 2D contour plot of $g$ and $f$ in Example 2
@@
Assume that $f,g$ are analytic functions on some open, connected domain $\Omega.$ Assume that $\{f,g\}=0$ and that the critical points of $f$ of even order are isolated. Then, there exists a continuous function $F:\mathbb{R}\rightarrow\mathbb{R}$ such that
@@
1.4.2. Properties of the set $\{\nabla\psi=0\}$ when $\{\psi,\Delta\psi\}=0$
Given the preceding discussion, the obstruction to having a global relation between $\Delta\psi$ and $\psi$ in a 2D Euler steady state is the existence of a curve of critical points of $\psi$. Since $\psi$ is analytic, we can consider the innermost such curve, $\Gamma$. In other words, in the open region enclosed by $\Gamma$ (possibly with the aid of a portion of the boundary $\partial\Omega$), called $\Omega_{\Gamma},$ $\psi$ has only isolated critical points. In $\Omega_{\Gamma}$, $\psi$ will satisfy a semilinear elliptic equation with nonlinearity $F$ and the full gradient of $\psi$ vanishes on $\Gamma$. Note that the curve $\Gamma$ may have singularities. In fact, if $\psi$ cannot be written as a solution of (1.4) globally, one of the following three cases must happen:
@@ 1.4.2. Properties of the set $\{\nabla\psi=0\}$ when $\{\psi,\Delta\psi\}=0$
    $\psi$ vanishes higher than quadratically on $\Gamma$;
@@
    $\psi$ vanishes only quadratically (somewhere) on $\Gamma$ and the associated nonlinearity $F$ is not analytic at $c=\psi(\Gamma)$;
@@
In both cases (i) and (ii), because $F$ is necessarily non-analytic at $c=\psi(\Gamma)$, it turns out that $\psi-c$ is necessarily single-signed in $\Omega_{\Gamma}$–this is a key aspect of the argument. In case (i), we can apply a known result of Brock [6] to deduce that $\psi$ must be radial; a nice feature of Brock’s result [6] is that it does not require the boundary of the domain to be smooth. The analysis in case (ii) is more difficult and will be discussed in more detail below. We give an illustration depicting the potential complexity of the level-set in Figure 3222The solid curves on the figure above are the set of critical points of $\psi$ in an elliptical domain. The boundary of the region $A$ demonstrates situation (i), where all points are critical points of degree higher than 2. The boundary of the region $B$ demonstrates the second case (ii), in the case where all points are quadratic. The regions $C$ and $D$ could represent the third case, where $F$ could be extended from $C$ to $D$, while $D$ falls into case (i) or (ii)..
Figure 3. Illustrations for a potential critical set of $\psi$
One snag in case (ii) is that we have to show that this curve $\Gamma$ is actually smooth and that $\psi$ is single-signed in its interior. For general analytic functions, it is well known that the zero set may be quite complicated and may contain multiple singular points of different types [34, 7], even more so in high dimensions. Note, for example, that the set of critical points in Example 2 in Section 1.4.1 has a singularity at the origin. In Figure 3, we have to show that case (ii) above actually corresponds to a region like $B$ in the figure (i.e., a region with a smooth boundary). A key observation here is that, when we have $\{\psi,\Delta\psi\}=0,$ the degree of vanishing of $\nabla\psi$ on each connected component of $\Gamma$ will necessarily be constant. This is established by using the constancy of $\Delta\psi$ in the levels of $\psi$. Once the constancy of the degree is established, we use an elementary fact: singularities in the zero set of an analytic function occur only at points where the degree of vanishing of the function changes. The constant degree of vanishing thus implies that $\Gamma$ is smooth.
In the case (iii), as $F$ is analytic in the boundary $\partial\Omega_{\Gamma}$, we can find an extension of $F$, such that $\psi$ would satisfy $\Delta\psi=F(\psi)$ in a domain containing $\Omega_{\Gamma}$. In practice, we do not work with case (iii) directly. Instead, based on the discussions on the case (i) and case (ii), we show if $\psi$ is neither radial nor a global solution of (1.4), then $\psi$ would have a nested family of curves of critical points. It then leads to a contradiction.
1.4.3. Moving Plane Arguments
For case (ii) of the previous subsection, we see that if a steady state $\psi$ does not satisfy the global relation $\Delta\psi=F(\psi)$, up to subtracting a constant, we end up with a simply connected domain $\Omega_{\Gamma}$ with analytic boundary where